# Random Feature Representation Boosting

### v1

@@ Abstract
We introduce Random Feature Representation Boosting (RFRBoost), a novel method for constructing deep residual random feature neural networks (RFNNs) using boosting theory. RFRBoost uses random features at each layer to learn the functional gradient of the network representation, enhancing performance while preserving the convex optimization benefits of RFNNs. In the case of MSE loss, we obtain closed-form solutions to greedy layer-wise boosting with random features. For general loss functions, we show that fitting random feature residual blocks reduces to solving a quadratically constrained least squares problem. We demonstrate, through numerical experiments on 91 tabular datasets for regression and classification, that RFRBoost significantly outperforms traditional RFNNs and end-to-end trained MLP ResNets, while offering substantial computational advantages and theoretical guarantees stemming from boosting theory.
Machine Learning, ICML, random features, residual neural networks, gradient boosting, gradient representation boosting, random feature neural networks
1Department of Mathematics, Imperial College London, UK
2School of Computer Science, University of St. Gallen, Switzerland
[…]
@@ 1 Introduction
as an Euler discretization of a dynamical system $d\Phi_{t}(x)=g_{t}(\Phi_{t})dt$ (E, 2017). Here $g_{t}$ represents a residual block at layer $t$, often expressed as $g_{t}(x)=A\sigma(Bx+b)$. This framework was later generalized to the setting of neural ODEs (Chen et al., 2018; Dupont et al., 2019; Kidger et al., 2020, 2021; Walker et al., 2024). The second point of view is that of gradient boosting, where a ResNet can be seen as an ensemble of weak, shallow neural networks of varying sizes, $\Phi_{T}(x)=\sum_{t=1}^{T}g_{t}(\Phi_{t-1}(x))$, derived by unravelling equation (1) (Veit et al., 2016). This led to the development of gradient representation boosting (Nitanda & Suzuki, 2018; Suggala et al., 2020), which studies residual blocks via functional gradients in the space of square integrable random vectors $L_{2}^{D}(\mu)$, where $\mu$ is the law of the data.
@@ 1.1 Contributions
  • Analytical Solutions and Algorithms: For MSE loss, we derive closed-form solutions to greedy layer-wise boosting using random features by solving what we term “sandwiched least squares problems”. For general losses, we show that fitting random feature residual blocks is equivalent to solving a quadratically constrained least squares problem.
@@ 1.2 Related literature
Random feature models: Random feature models use fixed, randomly generated features to map data into a high dimensional space, enabling efficient linear learning. These methods encompass a broad class of models studied under various names, including random Fourier features (Rahimi & Recht, 2007; Sriperumbudur & Szabo, 2015; Li et al., 2019; Kammonen et al., 2022; Davis et al., 2024), extreme learning machines (Huang et al., 2004, 2012; Huang, 2014), random features or RFNNs (Huang et al., 2006; Rahimi & Recht, 2008a, b; Rudi & Rosasco, 2017; Carratino et al., 2018; Yehudai & Shamir, 2019; Mei & Montanari, 2022; Bolager et al., 2023; Lanthaler & Nelsen, 2023; Ayme et al., 2024), reservoir computing (Jaeger, 2001; Lukoševičius & Jaeger, 2009; Gallicchio et al., 2017; Grigoryeva & Ortega, 2018; Tanaka et al., 2019; Hart et al., 2020; Gonon et al., 2023, 2024), and kernel methods (Kar & Karnick, 2012; Sinha & Duchi, 2016; Sun et al., 2018; Szabo & Sriperumbudur, 2019; Cheng et al., 2023; Wang et al., 2024; Wang & Feng, 2024). Recently, RFNNs and related approaches have demonstrated exceptional performance in both speed and accuracy across a wide variety of applications. These include solving partial differential equations (Nelsen & Stuart, 2021; Gonon, 2023; Gattiglio et al., 2024; Neufeld & Schmocker, 2024), online learning (Prabhu et al., 2024), time series classification (Dempster et al., 2020, 2023), and in mathematical finance (Jacquier & Zuric, 2023; Herrera et al., 2024). RFNNs have also been explored in the context of quantum computing (Innocenti et al., 2023; Gonon & Jacquier, 2023; Martínez-Peña & Ortega, 2023; Xiong et al., 2024), and in relation to random neural controlled differential equations and state-space models (Cuchiero et al., 2021; Cirone et al., 2023, 2024).
@@ 2 Functional Gradient Boosting
Let $(X,Y)\sim\mu$ be a random sample with its corresponding probability measure $\mu$. Denote by $X\sim\mu_{X}$ the features in $\mathbb{R}^{q}$, with targets $Y\sim\mu_{Y}$ in $\mathbb{R}^{d}$, and $\mu_{Y|X=x}$ the conditional law. Let $\widehat{\mu}=\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}$ be the empirical measure of $\mu$. We will be working in the Hilbert space $L_{2}^{D}(\mu)$, with inner product given by $\langle f,g\rangle_{L_{2}^{D}(\mu)}=\mathbb{E}_{\mu}[\langle f,g\rangle_{\mathbb{R}^{D}}]$. The concept of functional gradient plays a key role in gradient boosting:
@@ 2.1 Traditional Gradient Boosting
and is easily computed for empirical measures $\widehat{\mu}=\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}$ using the empirical risk ^$\widehat{\mathcal{R}}(F)$. The method for fitting $g_{t+1}$ varies between implementations and depends on the class of weak learners, as seen in modern approaches like XGBoost (Chen & Guestrin, 2016) which incorporates second-order information.
@@ 2.2 Gradient Representation Boosting
Unlike classical boosting, which boosts in label space, neural network gradient representation boosting aims to additively learn a feature representation by modelling $F$ via $F=F_{t}(X)=W_{t}^{\top}\Phi_{t}(X)$, where $\Phi_{t}=\sum_{s=0}^{t}\eta g_{s}$ is a gradient boosted feature representation and $W_{t}\in\mathbb{R}^{C\times d}$ is the top-level linear predictor. In other words, gradient representation boosting seeks to learn the feature representation $\Phi_{t}$ additively to build a single strong predictor, rather than constructing $F_{t}$ as an ensemble of weak predictors (Huang et al., 2018; Nitanda & Suzuki, 2018, 2020; Suggala et al., 2020). For a depth $t$ ResNet $\Phi_{t}$, defined recursively as
@@
However, in the context of random feature ResNets, using SGD for the joint optimization in Equation 3 undermines the computational benefits inherent to the random feature approach. As will be demonstrated in Section 3, by restricting $g_{t}$ to a simple residual block $g_{t}=\Delta_{t}f_{t}(\Phi_{t-1})$, with $f_{t}$ representing random features and $\Delta_{t}$ a linear map, we can derive closed-form analytical solutions for a two-stage approach to the optimization problem (3). This speeds up training and removes the need for hyperparameter tuning of the scale of the random features $f_{t}$ at each individual layer.
@@ 3 Random Feature Representation Boosting
$\Phi_{t}=\Phi_{t-1}+\eta\Delta_{t}f_{t},$
where $f_{t}=f_{t}(x,\Phi_{t-1}(x))\in\mathbb{R}^{p}$ are random features, $\Delta_{t}\in\mathbb{R}^{D\times p}$ is a linear map, $x$ is input data, and $\eta>0$ is the learning rate. We propose to use the theory of gradient representation boosting (Suggala et al., 2020) to derive optimal expressions for $\Delta_{t}$. We consider three different cases where $\Delta_{t}$ is either a scalar (learning optimal learning rate), a diagonal matrix (learning dimension-wise learning rate), or a dense matrix (learning the functional gradient).
@@ 3.2 Exact-Greedy Case for Mean Squared Error Loss
Recall that in the exact-greedy layer-wise paradigm, the objective at layer $t$ is to find a function $g_{t}\in\mathcal{G}_{t}$ that additively and greedily minimizes the empirical risk ^$\widehat{\mathcal{R}}(W_{t},\Phi_{t-1}+g_{t})$ for some linear map $W_{t}\in\mathcal{W}$, given a ResNet feature representation $\Phi_{t-1}$. We propose using functional gradient boosting to construct a residual block of the form $g_{t}=\Delta_{t}f_{t}$, where $\Delta_{t}\in\mathbb{R}^{D\times p}$ is a linear map, and $f_{t}$ is a random feature layer. We consider the cases where $\Delta_{t}$ is a scalar multiple of the identity, a diagonal matrix, or a general dense matrix. The procedure is outlined below:
@@
$\Delta_{t}$ $=\operatorname*{argmin}\limits_{\Delta}\widehat{\mathcal{R}}(W_{t-1},\Phi_{t-1}+\Delta f_{t}(x_{i}))$
$=\operatorname*{argmin}\limits_{\Delta}\frac{1}{n}\sum_{i=1}^{n}\|y_{i}-W_{t-1}^{\top}(\Phi_{t-1}(x_{i})+\Delta f_{t}(x_{i}))\|^{2}$
$=\operatorname*{argmin}\limits_{\Delta}\frac{1}{n}\sum_{i=1}^{n}\|r_{i}-W_{t-1}^{\top}\Delta f_{t}(x_{i})\|^{2},$ (5)
where $r_{i}=y_{i}-W_{t-1}^{\top}\Phi_{t-1}(x_{i})$ are the residuals of the model at layer $t-1$. We term problems of the form (3.2) sandwiched least squares problems, for which Theorem 3.1 provides the existence of closed-form analytical solutions which are fast to compute.
Step 2: After computing $\Delta_{t}$, we obtain the depth $t$ representation $\Phi_{t}$, and the top-level linear regressor $W_{t+1}$ is updated via multiple least squares
$\Phi_{t}=\Phi_{t-1}+\eta\Delta_{t}f_{t},$
@@
In practice, $\ell_{2}$ regularization is added to equations (3.2) and (6). Steps 1 and 2 are repeated for $T$ layers. The complete procedure is summarized in Algorithm 1.
@@
Let $r_{i}\in\mathbb{R}^{d},W\in\mathbb{R}^{D\times d},z_{i}\in\mathbb{R}^{p}$ for all $i\in[n]$. Let $\lambda>0$. Consider the setting of scalar $\Delta$, diagonal $\Delta$, and dense $\Delta\in\mathbb{R}^{D\times p}$. Write $Z\in\mathbb{R}^{N\times p}$ and $R\in\mathbb{R}^{N\times d}$ for the stacked data and residual matrices. Then the minimum of
$J(\Delta)$ $=\frac{1}{n}\sum_{i=1}^{n}\big{\|}r_{i}-W^{\top}\Delta z_{i}\big{\|}^{2}+\lambda\|\Delta\|_{F}^{2}$
@@
$\Delta_{\textnormal{scalar}}$ $=\frac{\langle R,ZW\rangle_{F}}{\|ZW\|_{F}^{2}+n\lambda},$
$\Delta_{\textnormal{diag}}=(WW^{\top}\odot Z^{\top}Z+\lambda I)^{-1}\textnormal{diag}(WR^{\top}Z),$
$\Delta_{\textnormal{dense}}=U\bigg{[}U^{\top}W$ $R^{\top}ZV\oslash\bigg{(}\lambda N\mathbf{1}$
@@
Algorithm 1 GreedyRFRBoost - MSE Loss
@@
     Solve sandwiched least squares $\Delta_{t}\leftarrow\operatorname*{argmin}\limits_{\Delta}\frac{1}{n}\sum_{i=1}^{n}\|r_{i}-W_{t-1}^{\top}\Delta f_{t,i}\|^{2}+\lambda\|\Delta\|_{F}^{2}$
     Build ResNet layer $\Phi_{t}\leftarrow\Phi_{t-1}+\eta\Delta_{t}f_{t}$
@@
Algorithm 2 GradientRFRBoost - General Loss
@@
     Fit gradient $\Delta_{t}\leftarrow least\_squares(f_{t},-\frac{\sqrt{n}G}{\|G\|_{F}},\lambda)$
     Find ‘amount of say’ with convex solver $\alpha_{t}\leftarrow\operatorname*{argmin}\limits_{\alpha\in\mathbb{R}}\frac{1}{n}\sum_{i=1}^{n}l\big{(}W^{\top}(\Phi_{t-1}+\alpha\Delta_{t}f_{t}),y_{i}\big{)}$
     Build ResNet layer $\Phi_{t}\leftarrow\Phi_{t-1}+\eta\alpha_{t}\Delta_{t}f_{t}$

### v2

@@ Abstract
We introduce Random Feature Representation Boosting (RFRBoost), a novel method for constructing deep residual random feature neural networks (RFNNs) using boosting theory. RFRBoost uses random features at each layer to learn the functional gradient of the network representation, enhancing performance while preserving the convex optimization benefits of RFNNs. In the case of MSE loss, we obtain closed-form solutions to greedy layer-wise boosting with random features. For general loss functions, we show that fitting random feature residual blocks reduces to solving a quadratically constrained least squares problem. Through extensive numerical experiments on tabular datasets for both regression and classification, we show that RFRBoost significantly outperforms RFNNs and end-to-end trained MLP ResNets in the small- to medium-scale regime where RFNNs are typically applied. Moreover, RFRBoost offers substantial computational benefits, and theoretical guarantees stemming from boosting theory.
Machine Learning, ICML, random features, gradient boosting, gradient representation boosting, residual neural networks, resnets, tabular data, supervised learning, random feature neural networks
@@ 1 Introduction
as an Euler discretization of a dynamical system $d\Phi_{t}(x)=g_{t}(\Phi_{t})dt$ (E, 2017). Here $g_{t}$ represents a residual block at layer $t$, often expressed as $g_{t}(x)=A\sigma(Bx+b)$. This framework was later generalized to the setting of neural ODEs (Chen et al., 2018; Dupont et al., 2019; Kidger et al., 2020, 2021; Walker et al., 2024). The second point of view is that of gradient boosting, where a ResNet can be seen as an ensemble of weak, shallow neural networks of varying sizes, $\Phi_{T}(x)=\sum_{t=1}^{T}g_{t}(\Phi_{t-1}(x))$, derived by unravelling equation (1) (Veit et al., 2016). This led to the development of gradient representation boosting (Nitanda & Suzuki, 2018; Suggala et al., 2020), which studies residual blocks via functional gradients in the space of square integrable random vectors $L_{2}^{D}(\mu)$, where $\mu$ is the distribution of the data.
@@ 1.1 Contributions
  • Analytical Solutions and Algorithms: For MSE loss, we derive closed-form solutions to greedy layer-wise boosting using random features by solving what we term “sandwiched least squares problems”, a special case of generalized Sylvester equations. For general losses, we show that fitting random feature residual blocks is equivalent to solving a quadratically constrained least squares problem.
@@ 1.2 Related literature
Random feature models: Random feature models use fixed, randomly generated features to map data into a high dimensional space, enabling efficient linear learning. These methods encompass a broad class of models studied under various names, including random Fourier features (Rahimi & Recht, 2007; Sriperumbudur & Szabo, 2015; Li et al., 2019; Kammonen et al., 2022; Davis et al., 2024), extreme learning machines (Huang et al., 2004, 2012; Huang, 2014), random features or RFNNs (Huang et al., 2006; Rahimi & Recht, 2008a, b; Rudi & Rosasco, 2017; Carratino et al., 2018; Yehudai & Shamir, 2019; Mei & Montanari, 2022; Bolager et al., 2023; Lanthaler & Nelsen, 2023; Ayme et al., 2024), reservoir computing (Jaeger, 2001; Lukoševičius & Jaeger, 2009; Gallicchio et al., 2017; Grigoryeva & Ortega, 2018; Tanaka et al., 2019; Hart et al., 2020; Gonon et al., 2023, 2024), kernel methods (Kar & Karnick, 2012; Sinha & Duchi, 2016; Sun et al., 2018; Szabo & Sriperumbudur, 2019; Cheng et al., 2023; Wang et al., 2024; Wang & Feng, 2024), and scattering networks (Bruna & Mallat, 2013; Cotter & Kingsbury, 2017; Oyallon et al., 2019; Trockman et al., 2023). Recently, RFNNs and related approaches have demonstrated exceptional performance in both speed and accuracy across a wide variety of applications. These include solving partial differential equations (Nelsen & Stuart, 2021; Gonon, 2023; Gattiglio et al., 2024; Neufeld & Schmocker, 2024), online learning (Prabhu et al., 2024), time series classification (Dempster et al., 2020, 2023), and in mathematical finance (Jacquier & Zuric, 2023; Herrera et al., 2024). RFNNs have also been explored in the context of quantum computing (Innocenti et al., 2023; Gonon & Jacquier, 2023; Martínez-Peña & Ortega, 2023; Xiong et al., 2024), and in relation to random neural controlled differential equations and state-space models (Cuchiero et al., 2021; Cirone et al., 2023, 2024).
@@ 2 Functional Gradient Boosting
Let $(X,Y)\sim\mu$ be a random sample with its corresponding probability measure $\mu$. Denote by $X\sim\mu_{X}$ the features in $\mathbb{R}^{q}$, with targets $Y\sim\mu_{Y}$ in $\mathbb{R}^{d}$, and $\mu_{Y|X=x}$ the conditional law. Let $\widehat{\mu}=\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}$ be the empirical measure of $\mu$. We will work in the Hilbert space $L_{2}^{D}(\mu)$, with inner product given by $\langle f,g\rangle_{L_{2}^{D}(\mu)}=\mathbb{E}_{\mu}[\langle f,g\rangle_{\mathbb{R}^{D}}]$. The concept of functional gradient plays a key role in gradient boosting:
@@ 2.1 Traditional Gradient Boosting
and is easily computed for empirical measures $\widehat{\mu}=\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}$ using the empirical risk ^$\widehat{\mathcal{R}}(F)$. The method for fitting $g_{t+1}$ varies between implementations and depends on the class of weak learners, as seen in modern approaches like XGBoost (Chen & Guestrin, 2016) which incorporates second-order information.
@@ 2.2 Gradient Representation Boosting
Unlike classical boosting, which boosts in label space, neural network gradient representation boosting aims to additively learn a feature representation by modelling $F$ via $F=F_{t}(X)=W_{t}^{\top}\Phi_{t}(X)$, where $\Phi_{t}=\sum_{s=0}^{t}\eta g_{s}$ is a gradient boosted feature representation and $W_{t}\in\mathbb{R}^{D\times d}$ is the top-level linear predictor. In other words, gradient representation boosting seeks to learn the feature representation $\Phi_{t}$ additively to build a single strong predictor, rather than constructing $F_{t}$ as an ensemble of weak predictors (Huang et al., 2018; Nitanda & Suzuki, 2018, 2020; Suggala et al., 2020). For a depth $t$ ResNet $\Phi_{t}$, defined recursively as
@@
However, in the context of random feature ResNets, using SGD for the joint optimization in Equation 3 undermines the computational benefits inherent to the random feature approach. As will be demonstrated in Section 3, by restricting $g_{t}$ to a simple residual block $g_{t}=A_{t}f_{t}(\Phi_{t-1})$, with $f_{t}$ representing random features and $A_{t}$ a linear map, we can derive closed-form analytical solutions for a two-stage approach to the optimization problem (3). This speeds up training and removes the need for hyperparameter tuning of the scale of the random features $f_{t}$ at each individual layer.
@@ 3 Random Feature Representation Boosting
$\Phi_{t}=\Phi_{t-1}+\eta A_{t}f_{t},$
where $f_{t}=f_{t}(x,\Phi_{t-1}(x))\in\mathbb{R}^{p}$ are random features, $A_{t}\in\mathbb{R}^{D\times p}$ is a linear map, $x$ is input data, and $\eta>0$ is the learning rate. We propose to use the theory of gradient representation boosting (Suggala et al., 2020) to derive optimal expressions for $A_{t}$. We consider three different cases where $A_{t}$ is either a scalar (learning optimal learning rate), a diagonal matrix (learning dimension-wise learning rate), or a dense matrix (learning the functional gradient). Note that the scalar and diagonal cases assume that $p=D$.
@@ 3.2 Exact-Greedy Case for Mean Squared Error Loss
Recall that in the exact-greedy layer-wise paradigm, the objective at layer $t$ is to find a function $g_{t}\in\mathcal{G}_{t}$ that additively and greedily minimizes the empirical risk ^$\widehat{\mathcal{R}}(W_{t},\Phi_{t-1}+g_{t})$ for some linear map $W_{t}\in\mathcal{W}$, given a ResNet feature representation $\Phi_{t-1}$. We propose using functional gradient boosting to construct a residual block of the form $g_{t}=A_{t}f_{t}$, where $A_{t}\in\mathbb{R}^{D\times p}$ is a linear map, and $f_{t}$ is a random feature layer. We consider the cases where $A_{t}$ is a scalar multiple of the identity, a diagonal matrix, or a general dense matrix. The procedure is outlined below:
@@
$A_{t}$ $=\operatorname*{argmin}\limits_{A}\widehat{\mathcal{R}}(W_{t-1},\Phi_{t-1}+Af_{t}(x_{i}))$
$=\operatorname*{argmin}\limits_{A}\frac{1}{n}\sum_{i=1}^{n}\|y_{i}-W_{t-1}^{\top}(\Phi_{t-1}(x_{i})+Af_{t}(x_{i}))\|^{2}$
$=\operatorname*{argmin}\limits_{A}\frac{1}{n}\sum_{i=1}^{n}\|r_{i}-W_{t-1}^{\top}Af_{t}(x_{i})\|^{2},$ (5)
where $r_{i}=y_{i}-W_{t-1}^{\top}\Phi_{t-1}(x_{i})$ are the residuals of the model at layer $t-1$. We term problems of the form (3.2) sandwiched least squares problems, for which Theorem 3.1 below provides the existence of closed-form analytical solutions which are fast to compute.
Step 2: After computing $A_{t}$, we obtain the depth $t$ representation $\Phi_{t}=\Phi_{t-1}+\eta A_{t}f_{t}$. The top-level linear regressor $W_{t+1}$ is then updated via multiple least squares:
@@
In practice, $\ell_{2}$ regularization is added to equations (3.2) and (6). Steps 1 and 2 are repeated for $T$ layers. The complete procedure is detailed in Algorithm 1.
@@
Let $r_{i}\in\mathbb{R}^{d},W\in\mathbb{R}^{D\times d},z_{i}\in\mathbb{R}^{p}$ for all $i\in[n]$. Let $\lambda>0$. Consider the setting of scalar $A$, diagonal $A$, and dense $A\in\mathbb{R}^{D\times p}$. Write $Z\in\mathbb{R}^{N\times p}$ and $R\in\mathbb{R}^{N\times d}$ for the stacked data and residual matrices. Then the minimum of
$J(A)$ $=\frac{1}{n}\sum_{i=1}^{n}\big{\|}r_{i}-W^{\top}Az_{i}\big{\|}^{2}+\lambda\|A\|_{F}^{2}$
@@
$A_{\textnormal{scalar}}$ $=\frac{\langle R,ZW\rangle_{F}}{\|ZW\|_{F}^{2}+n\lambda},$
$A_{\textnormal{diag}}=(WW^{\top}\odot Z^{\top}Z+\lambda I)^{-1}\textnormal{diag}(WR^{\top}Z),$
$A_{\textnormal{dense}}=U\bigg{[}U^{\top}W$ $R^{\top}ZV\oslash\bigg{(}\lambda N\mathbf{1}$
@@
Algorithm 1 Greedy RFRBoost — MSE Loss
@@
     Solve sandwiched least squares $A_{t}\leftarrow\operatorname*{argmin}\limits_{A}\frac{1}{n}\sum_{i=1}^{n}\|r_{i}-W_{t-1}^{\top}Af_{t,i}\|^{2}+\lambda\|A\|_{F}^{2}$
     Build ResNet layer $\Phi_{t}\leftarrow\Phi_{t-1}+\eta A_{t}f_{t}$
@@
Algorithm 2 Gradient RFRBoost — General Loss
@@
     Fit gradient $A_{t}\leftarrow least\_squares(f_{t},-\frac{\sqrt{n}G}{\|G\|_{F}},\lambda)$
     Solve line search with convex solver $\alpha_{t}\leftarrow\operatorname*{argmin}\limits_{\alpha\in\mathbb{R}}\frac{1}{n}\sum_{i=1}^{n}l\big{(}W^{\top}(\Phi_{t-1}+\alpha A_{t}f_{t}),y_{i}\big{)}$
     Build ResNet layer $\Phi_{t}\leftarrow\Phi_{t-1}+\eta\alpha_{t}A_{t}f_{t}$